# appAdditional References Appendix Ensuring Medical AI Safety: Explainable AI-Driven Detection and Mitigation of Spurious Model Behavior and Associated Data

### v1

@@ 7 Limitations
While all steps in our framework are semi-automated, they require human supervision from domain experts, e.g., to validate outlier concepts, inspect detected bias samples, and determine which concepts should be unlearned. Below, we discuss additional challenges associated with each step of our framework.
@@
Bias Modeling: Without explicit concept disentanglement (e.g., via SAEs), concept representations may be non-orthogonal, leading to entangled concepts that negatively impact data annotation and bias mitigation. An additional limitation poses the linear separability assumption of CAVs, as there is no guarantee for the existence of a layer perfectly separating concept samples. This makes the layer selection an additional hyperparameter, and a suboptimal choice can exacerbate issues like concept entanglement. Non-localizable artifacts (e.g., color shifts) are another challenge, as they cannot be effectively modeled using spatial masks in input space. Lastly, poorly chosen data samples representing the concept can limit the accuracy of bias representations.
@@ 8 Conclusions
In this work, we introduce an XAI-based framework for the identification of spurious shortcut behavior in DNNs from both data and model perspective. We utilize concept-based bias representations for the semi-automated computation of sample- and pixel level bias annotations. Our work is integrated into the Reveal2Revise framework, providing valuable insights for bias mitigation and re-evaluation. We successfully demonstrated the applicability of the framework by identifying and mitigation spurious correlations caused by controlled and real-world data artifacts in four medical datasets across two modalities, using VGG16, ResNet50, and ViT model architectures. Future work may explore the identification and mitigation of biases in disentangled concept spaces, e.g., leveraging SAEs. Another promising direction is the integration of expected concepts for a more targeted search for unexpected concepts.
@@ Appendix A Appendix
In the following, we will provide additional details and mathematical foundations for CAV-based bias modeling approaches in Sec. A.1, biased sample retrieval in Sec. A.2, and bias mitigation approaches in Sec. A.3. We will then provide experimental details regarding the datasets in Sec. A.4 and model training in Sec. A.5. This is followed by additional experimental results in Sec. A.6, including XAI-driven shortcut identification for both ECG and vision data (Sec. A.6.1), biased sample retrieval (Sec. A.6.2), spatial bias localization (Sec. A.6.3), and bias mitigation (Sec. A.6.4).
A.1 Bias Modeling
@@
A.2 Biased Sample Retrieval
@@
with latent relevances w.r.t. class computed on layer by a local attribution method, e.g., using LRP. As the relevance is computed in a class-specific manner, this approach further allows distinguishing concepts that are artifactual for certain classes but valid features for others.
A.3 Bias Mitigation
A.3.1 Regularization-based Bias Mitigation
@@
Bias mitigation approaches on input level require sample-wise pixel-level annotations, indicating the presence of a data artifact, available as prior knowledge. This prior information can be used to align expected behavior with the model’s prediction stategy. Specifically, RRR [63] introduces a loss term penalizing the alignment between the input gradient, i.e., the gradient of the prediction w.r.t. the input, and (binary) input masks localizing the data artifact. Hence, the model is penalized for paying attention on undesired regions. Given an input sample along with bias localization mask , the loss for RRR is defined as
@@
To overcome aforementioned limitations, recent work introduces the ClArC framework [3] that leverages latent concept representations using CAVs for bias mitigation which has two advantages. First, using the latent space alleviates the reliance on spatial locations and therefore allows the correction of unlocalized biases, such as color shifts. Moreover, the annotation costs are substantially lower, since only sample-level concept labels are required for the computation of CAVs, instead of pixel-level artifact masks. Given the bias representation in the form of a CAV, the ClArC framework unlearns related model behavior by either adding or removing activations in the artifact direction. Built upon this idea, Dreyer et al. [28] introduce RR-ClArC, utilizing a loss function inspired by RRR which penalizes the feature use measured through the latent gradient pointing into the direction of the bias, as modeled via the CAV. More precisely, given a bias direction , the latent right reason loss is defined as
@@
A.3.2 Post-hoc Model Editing
@@
A.4 Dataset Details
We run experiments for the datasets ISIC2019 for image-based melanoma detection, HyperKvasir for the identification of abnormal conditions in the gastrointestinal track, CheXpert with chest radiographs, and PTB-XL for the prediction of cardiovascular diseases from ECG data. ISIC2019 consists of 25,331 samples classified into eight lesion types, both malignant and benign. The HyperKvasir dataset contains 10,662 samples labeled into 23 classes of findings. We aggregate the findings into two classes representing findings with and without disease to consider the task as a binary classification task. The CheXpert dataset is a collection of 224,316 chest radiographs of 65,240 patients with labels for 14 conditions, including cardiomegaly. We solve a binary classification task by predicting the presence of cardiomegaly. We use a subset of 28,878 samples with additional annotations for the presence of pacemakers provided by Weng et al. [75]. Lastly, in addition to vision data, we use the PTB-XL dataset containing 21,837 records of 10 second 12-lead ECG data (time series data) and labels for 23 cardiovascular conditions. We report details for all considered datasets in Tab. 2. Specifically, we report the number of classes and samples, train/validation/test-splits and considered data artifacts. We use ISIC2018 for multi-class classification and consider the real-world artifacts ruler, skin-marker, and reflection, as well as the artificial brightness artifact inserted into of samples of class “melanoma”. For ruler (all classes) and band-aid (only occurring in “melanocytic nevus”), we use labels indicating the existence of the artifact provided by Anders et al. [3]. For HyperKvasir, we split the classes into “normal”/“no-disease” and “abnormal”/“disease“ to implement the task as a binary classification task, as described in Tab. 3. We insert a timestamp artifact into of the the training samples with label “disease”, and, in addition, consider the real-world artifact insertion tube. The latter was primarily detected in samples from class “no disease”. For CheXpert, we solve a binary classification task to predict the presence of the class “cardiomegaly”. We limit our experiments to a subset of the CheXpert dataset for which Weng et al. [75] provide labels for the existence of the naturally occuring pacemaker artifact (mainly in class “cardiomegaly”). In addition, we artificially increase the brightness in of “cardiomegaly” samples as a controlled artifact. Lastly, we use the PTB-XL dataset with ECG-data for multi-class classification into cardiovascular conditions. As a controlled artifact, we insert a static noise (constant high value) into the first second of one lead (II-lead) for of samples of class “left ventricular hypertrophy”.
@@ A.6 Additional Results
A.6.1 Detection of Spurious Model Behavior
@@
A.6.2 Biased Sample Retrieval
@@
A.6.3 Spatial Bias Localization

### v2

@@ 7 Limitations
While all steps in the extended Reveal2Revise framework are semi-automated, they require human supervision from domain experts, e.g., to validate outlier concepts, inspect detected bias samples, and determine which concepts should be unlearned. Below, we discuss additional challenges associated with each step of the framework.
@@
Bias Modeling: Without explicit concept disentanglement (e.g., via SAEs), concept representations may be non-orthogonal, leading to entangled concepts that negatively impact data annotation and bias mitigation. Additionally, the linear separability assumption of CAVs poses limitations, as there is no guarantee for the existence of a layer perfectly separating concept samples. This makes the layer selection an additional hyperparameter, and a suboptimal choice can exacerbate issues like concept entanglement. Non-localizable artifacts (e.g., color shifts) are another challenge, as they cannot be effectively modeled using spatial masks in input space. Lastly, poorly chosen data samples representing the concept can limit the accuracy of bias representations.
@@ 8 Conclusions
In this work, we reviewed the steps of Reveal2Revise, a comprehensive interpretability-based framework for the detection and mitigation of spurious shortcut behavior in DNNs, and enhanced the framework with bias annotation capabilities. Specifically, we utilized concept-based bias representations for the semi-automated computation of sample- and feature-level bias annotations, providing valuable insights for the bias mitigation and re-evaluation steps. We successfully demonstrated the applicability of the extended Reveal2Revise framework by identifying and mitigating spurious correlations caused by controlled and real-world data artifacts in four medical datasets across two modalities, using VGG16, ResNet50, and ViT model architectures. Future work may explore the identification and mitigation of biases in disentangled concept spaces, e.g., leveraging SAEs. Another promising direction is the integration of expected concepts for a more targeted search for unexpected concepts.
@@ Appendix A Appendix
In the following, we will provide an algorithmic overview of the extended Reveal2Revise framework in Sec. A.1, additional details and mathematical foundations for CAV-based bias modeling approaches in Sec. A.2, biased sample retrieval in Sec. A.3, and bias mitigation approaches in Sec. A.4. We will then provide experimental details regarding the datasets in Sec. A.5 and model training in Sec. A.6. This is followed by additional experimental results in Sec. A.7, including XAI-driven shortcut identification for both ECG and vision data (Sec. A.7.1), biased sample retrieval (Sec. A.7.2), spatial bias localization (Sec. A.7.3), and bias mitigation (Sec. A.7.4).
A.1 Reveal2Revise Framework
In addition to the visualization in Fig. 1 (left), we provide an algorithmic overview of the Reveal2Revise framework in Algorithm 1. It includes the steps (a) reveal to detect model biases, (b) bias modeling to obtain a suitable representation of the bias, (c) revise to correct the model behavior with a bias mitigation approach, and (d) evaluate to measure if the spurious model behavior was unlearned successfully.
In addition, we provide an overivew of our proposed extension to the (a) reveal and (b) bias modeling steps in Algorithm 2. Specifically, the application of bias identification approaches either from the model or data perspective provides an initial set of bias-affected samples, which can be utilized to construct the bias concept labels ( ). The bias labels are leveraged to train a CAV as bias representation ( ), which is subsequently employed to identify additional biased samples ( ). Specifically, as conceptualized in Fig. 3, this process involves calculating bias scores (see Eq. 1) for samples that are currently labeled as not containing the bias according to the current labels . The top high-scoring samples are subject to manual inspection to further refine the bias labels . The process is repeated until convergence ( ). Finally, the updated CAV is utilized to spatially localize the bias in affected samples ( ). The algorithm outputs refined bias labels , CAV as bias representation, and spatial localization masks .
A.2 Bias Modeling
@@
A.3 Biased Sample Retrieval
@@
with latent relevances for sample w.r.t. class computed on layer by a local attribution method, e.g., using LRP. As the relevance is computed in a class-specific manner, this approach further allows distinguishing concepts that are artifactual for certain classes but valid features for others.
A.4 Bias Mitigation
A.4.1 Regularization-based Bias Mitigation
@@
Bias mitigation approaches on input level require sample-wise pixel-level annotations, indicating the presence of a data artifact, available as prior knowledge. This prior information can be used to align expected behavior with the model’s prediction stategy. Specifically, RRR [63] introduces a loss term penalizing the alignment between the input gradient, i.e., the gradient of the prediction w.r.t. the input, and (binary) input masks localizing the data artifact. Hence, the model is penalized for paying attention on undesired regions. Given an input sample along with bias localization mask , the RRR loss term is defined as follows:
@@
To overcome aforementioned limitations, recent work introduces the ClArC framework [3] that leverages latent concept representations using CAVs for bias mitigation which has two advantages. First, using the latent space alleviates the reliance on spatial locations and therefore allows the correction of unlocalized biases, such as color shifts. Moreover, the annotation costs are substantially lower, since only sample-level concept labels are required for the computation of CAVs, instead of pixel-level artifact masks. Given the bias representation in the form of a CAV, the ClArC framework unlearns related model behavior by either adding or removing activations in the artifact direction. Built upon this idea, Dreyer et al. [29] introduce RR-ClArC, utilizing a loss function inspired by RRR which penalizes the feature use measured through the latent gradient pointing into the direction of the bias, as modeled via the CAV. More precisely, given a bias direction , the latent right reason loss is defined as follows:
@@
A.4.2 Post-hoc Model Editing
@@
A.5 Dataset Details
Our experiments include the following datasets: ISIC2019 for image-based melanoma detection, HyperKvasir for the identification of abnormal conditions in the gastrointestinal track, CheXpert with chest radiographs, and PTB-XL for the prediction of cardiovascular diseases from ECG data. ISIC2019 consists of 25,331 samples classified into eight lesion types, both malignant and benign. The HyperKvasir dataset contains 10,662 samples labeled into 23 classes of findings. We aggregate the findings into two classes representing findings with and without disease to consider the task as a binary classification task. The CheXpert dataset is a collection of 224,316 chest radiographs of 65,240 patients with labels for 14 conditions, including cardiomegaly. We use a subset of 28,878 samples with additional annotations for the presence of pacemakers provided by Weng et al. [76]. Lastly, in addition to vision data, we use the PTB-XL dataset containing 21,837 records of 10 second 12-lead ECG data (time series data) and labels for 23 cardiovascular conditions. We report details for all considered datasets in Tab. 2. Specifically, we report the number of classes and samples, train/validation/test-splits and considered data artifacts.
We use ISIC2019 for multi-class classification and consider the real-world artifacts ruler, skin-marker, and reflection, as well as the artificial brightness artifact inserted into of samples of class “melanoma”. For ruler (all classes) and band-aid (only occurring in “melanocytic nevus”), we use labels indicating the existence of the artifact provided by Anders et al. [3]. For HyperKvasir, we split the classes into “normal”/“no-disease” and “abnormal”/“disease“ to implement the task as a binary classification task, as described in Tab. 3. We insert a timestamp artifact into of the the training samples with label “disease”, and, in addition, consider the real-world artifact insertion tube. The latter was primarily detected in samples from class “no disease”. For CheXpert, we solve a binary classification task to predict the presence of the class “cardiomegaly”. We consider the pacemaker artifact dominantly occurring in class “cardiomegaly”, with artifact labels provided by Weng et al. [76]. In addition, we artificially increase the brightness in of “cardiomegaly” samples as a controlled artifact. Lastly, we use the PTB-XL dataset with ECG-data for multi-class classification into cardiovascular conditions. As a controlled artifact, we insert a static noise (constant high value) into the first second of one lead (II-lead) for of samples of class “left ventricular hypertrophy”.
@@ A.7 Additional Results
A.7.1 Detection of Spurious Model Behavior
@@
A.7.2 Biased Sample Retrieval
@@
A.7.3 Spatial Bias Localization